Mathematics 132 – Calculus for Physical and Life Sciences 2 Makeup Quiz – Wednesday March 12, 2008 Your Name: Some possibly useful formulas:

sin^{2} θ + cos^{2} θ = 1,  tan^{2} θ + 1 = sec^{2} θ,  sin(2θ) = 2 sin θ cos θ.

Do all work on this sheet (front and back).

(5) A) Integrate:

∫ x^{3}\sqrt{16 − x^{2}} dx

(5) B) Integrate:

∫ \frac{x^{4} + 3x + 5}{x^{3} + x^{2}} dx